OpenStudy (anonymous):

Defferentiate y = e^4x cosx

OpenStudy (anonymous):

e^4xcosx * (4cosx - 4xsinx)

OpenStudy (anonymous):

Answer is -e^4x sinx + 4e^4x cosx but i'm still confused.

OpenStudy (anonymous):

but y r u confused

OpenStudy (anonymous):

okay i think i get it better now. so, u = e^ 4x u' = 4e^4x, v = cosx, v' = -sinx then I use u'v +v'u?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Does it matter if i have it backwards? like 4e^4x cos-e^4x sinx instead of -e^4x sinx + 4e^4x cosx?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay. Thanks!

OpenStudy (anonymous):

mention not